The client is thinking of spending $32,500 in direct mail next week. What is your best estimate of what the following week’s sales will be at that level of spending?

Week Sales Direct Mail Corrected
1 $121,230 9800 42047
2 $212,090 21200 9800
3 $99,980 561000 21200
4 $429,780 41300 56100
5 $496,370 26700 41300
6 $316,450 32100 26700
7 $286,980 52900 32100
8 $496,110 73100 52900
9 $389,080 69500 73100
10 $787,350 54400 69500
11 $446,310 8700 54400
12 $389,410 29900 8700
13 $420,040 24300 29900
14 $629,380 42300 24300
15 $419,370 82000 42300
16 $740,070 60200 82000
17 $498,730 48900 60200
18 $621,780 27900 48900
19 $317,620 37600 27900
20 $427,270 68621 37600
Answer #1

using excel>data analysis>regression

we have

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.367221
R Square 0.134851
Adjusted R Square 0.086788
Standard Error 171962.5
Observations 20
ANOVA
df SS MS F Significance F
Regression 1 8.3E+10 8.3E+10 2.805675 0.111216
Residual 18 5.32E+11 2.96E+10
Total 19 6.15E+11
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 465778.1 44800.48 10.39672 4.89E-09 371655.8 559900.5
Direct Mail -0.56117 0.335024 -1.67501 0.111216 -1.26503 0.142689

the regression equation is

Sales = 465778.1 -0.56117 * Direct Mail

for direct mail = 32,500

estimated sale = 465778.1 -0.56117*32500 = 447540.075